OpenStudy (anonymous):

For positive real numbers \(\Large a, b, c\), find the minimum integer value possible of the following equation: \(\Large 6a^3+9b^3+32c^3+\frac{1}{4abc}\)

OpenStudy (kainui):

Hmm, is that last positive sign supposed to be an equals sign, or is there more here?

OpenStudy (aum):

An equation has to have an equal sign which is missing here.

OpenStudy (aum):

Is it supposed to be "the following expression" instead of "the following equation" ?

OpenStudy (anonymous):

@aum yes Sorry I write wrong

OpenStudy (anonymous):

Hint: Using the AM-GM Inequality

OpenStudy (anonymous):

The main key in solving this equation is using the AM-GM Inequality. -Using the AM-GM Inequality with \(\Large x_1=6a^3, x_2=9b^3, x_3=32c^3\) we obtain, \(\Large \frac{6a^3+9b^3+32c^3}{3} \ge \sqrt[3]{1728a^3b^3c^3}\) Which simplifies nicely into, \(\Large 6a^3+9b^3+32c^3 \ge 36abc\) Thus, adding \(\Large \frac{1}{4abc}\) on both sides we get, \(\Large 6a^3+9b^3+32c^3+ \frac{1}{4abc} \ge 36abc+\frac{1}{4abc}\) -Applying AM-GM on RHS on the inequality, \(\Large x_1=36abc, x_2=\frac{1}{4abc}\), we attain, \(\huge \frac{36abc+ \frac{1}{4abc}}{2}\ge \sqrt[2]{36abc\times \frac{1}{4abc}}\\\huge 36abc+\frac{1}{4abc}\ge 6\) Thus, \(\Large 6a^3+9b^3+32c^3+ \frac{1}{4abc} \ge 36abc+\frac{1}{4abc}\ge 6\) Minimum value is 6.

OpenStudy (aum):

Nice problem!

OpenStudy (anonymous):

:) thanks